# Grammar-Based Code Representation: Is It a Worthy Pursuit for LLMs?

### v1

@@ Abstract
Grammar serves as a cornerstone in programming languages and software engineering, providing frameworks to define the syntactic space and program structure. Existing research demonstrates the effectiveness of grammar-based code representations in small-scale models, showing their ability to reduce syntax errors and enhance performance. However, as language models scale to the billion level or beyond, syntax-level errors become rare, making it unclear whether grammar information still provides performance benefits. To explore this, we develop a series of billion-scale GrammarCoder models, incorporating grammar rules in the code generation process. Experiments on HumanEval (+) and MBPP (+) demonstrate a notable improvement in code generation accuracy. Further analysis shows that grammar-based representations enhance LLMs’ ability to discern subtle code differences, reducing semantic errors caused by minor variations. These findings suggest that grammar-based code representations remain valuable even in billion-scale models, not only by maintaining syntax correctness but also by improving semantic differentiation.111 https://github.com/LIANGQINGYUAN/GrammarCoder.
@@ 1 Introduction
Context-free grammars are the fundamental way to specify the syntactic space of a programming language, and with the grammar specified, a program can be parsed into a syntax tree, revealing its structure Aho et al. (1986). Building on this foundation, leveraging grammatical knowledge (e.g., grammar rules) to pre-train large language models (LLMs) has emerged as a promising strategy for code-related tasks, such as code generation Zhu et al. (2024); Sun et al. (2020); Guo et al. (2020).
@@
Existing research has explored grammar-based code representation Jiang et al. (2021); Guo et al. (2022); Wang et al. (2021a); Zhu et al. (2024); Sun et al. (2020); Xiong and Wang (2022); Rabinovich et al. (2017), where each grammar rule serves as an identity token, and a sequence of grammar rules and terminal tokens represents the program. Figure 1 illustrates a program that determines whether the sum of two integers is odd (top left), along with its corresponding abstract syntax tree (AST) representation (right) and grammar-based representation (bottom left). The grammar-based representation is derived by performing a preorder traversal on the AST. Each grammar rule is extracted independently (e.g., module $\rightarrow$ function_definition’), while terminals are tokenized using a standard tokenizer (e.g., get’). Grammar-based representation has been shown to be effective in preventing syntax errors in encoder-decoder architecture Zhu et al. (2024). Moreover, it facilitates program analysis and enables the pruning of incorrect branches (e.g., filtering out type-error programs Xiong and Wang (2022); Zhu et al. (2023)) during code generation, thereby enhancing accuracy. Due to these benefits, many code generation models adopt grammar-based representation Sun et al. (2019); Zhu et al. (2024).
@@
However, as language models scale to the billion-parameter level and beyond, extensive pre-training on large code datasets enables them to implicitly learn syntax rules, making syntax errors increasingly rare OpenAI (2024); Yang et al. (2024); Team (2024); DeepSeek-AI et al. (2024). For example, even 1B scale models, such as DeepSeek-Coder Guo et al. (2024) and Qwen2.5 Team (2024), achieve high accuracy in code generation, consistently producing syntactically valid code. This phenomenon suggests that large models are able to understand the structure of the program and raises a critical question: Is grammar-based code representation still beneficial for billion-scale LLMs?
@@
To answer this question, we conduct an experiment comparing grammar-based representation and token-based representation approaches on 1.3B and 1.5B parameter models, respectively. The results demonstrate that grammar-based models (i.e., GrammarCoder-1.3B-Base and GrammarCoder-1.5B-Base) significantly outperform token-based models, even though token-based models rarely make syntax errors. For example, on the MBPP dataset Austin et al. (2021), GrammarCoder-1.3B-Base achieves an almost seven percentage point improvement in Pass@1 compared to DeepSeek-Coder-1.3B-Base trained on the same data. This suggests that grammar rules enhance code generation beyond syntax correction, even in billion-scale models.
@@
  • We release a series of code LLMs trained with grammar rules, providing a valuable resource for further research.
@@ 2.1 Model Overview
We propose GrammarCoder, a grammar-based model built on a decoder-only architecture Vaswani et al. (2017); Radford et al. (2018), which excels in auto-regressive tasks like code generation, completion, and translation OpenAI (2024); Guo et al. (2024); Team (2024); Hui et al. (2024); DeepSeek-AI et al. (2024). To enhance its ability of code generation, we apply continued pre-training and instruction tuning on existing code model weights (i.e., DeepSeek-Coder-1.3B-Base and Qwen2.5-1.5B-Base), expanding its knowledge base. A.1 provides the configuration of the base model we used. In this section, we first introduce our grammar-based code representation. Then, we describe our training strategy and corpus.
@@ 2.4 Training Corpus
We organize our training corpus in two stages: base model training and instruction tuning. Python is selected as the primary programming language for data collection due to its rich syntax and widespread use in diverse programming paradigms. This makes it an ideal candidate for evaluating the effectiveness of grammar-based representations.
For base model training, we sample 10B tokens of Python code from TheStackV2 Lozhkov et al. (2024) dataset as the primary training data. Additionally, inspired by previous studies Huang et al. (2024), we sample 0.5B tokens of self-contained code textbooks from open-source datasets Huang et al. (2024); Nakamura et al. (2025) to enhance the model’s adaptability to real-world interactive scenarios, bridging the gap between standard pre-training and practical applications.
For instruction tuning, we leverage publicly available instruction datasets Huang et al. (2024); Nakamura et al. (2025) and employ the data synthesis Wei et al. (2024a, b) approach to collect a total of 6B tokens of instruction data. This ensures the model is better aligned with instruction-following tasks, improving its ability to handle real-world programming scenarios. A.3 provides detailed information about the training datasets.
@@
GrammarCoder-1.5B-Base 63.4 (59.1) 64.8 (55.3)
Table 1: Comparison of code generation performance between token-based and grammar-based models. The CPT refers to continued pre-training, while the SFT denotes supervised fine-tuning.
@@ 3 Experiments
To evaluate the performance of grammar-based code representations, we develop two sets of models, one with grammar-based code representation and one with token-based code representation. These models are built through continued pre-training from open-source code models, DeepSeek-Coder-1.3B-Base and Qwen2.5-1.5B, on high-quality code data. We begin by evaluating these models on code generation tasks, which are among the most widely recognized and commonly used benchmarks for assessing code-related capabilities (i.e., Experiment I in 3.1).
To further explore the differences between grammar-based and token-based representations, we analyze the reason contributing to the performance gains of grammar-based representation (i.e., Experiment II in 3.2).
@@
DeepSeek-Coder-1.3B-Base Guo et al. (2024) 34.8 28.7 56.7 47.9
Qwen2.5-1.5B-Base Team (2024) 37.2 32.9 60.2 49.6
OpenCoder-1.5B-Base Huang et al. (2024) 54.3 49.4 70.6 58.7
Yi-Coder-1.5B AI et al. (2025) 41.5 32.9 27.0 22.2
CodeGemma-2B-Base Team et al. (2024) 26.8 20.7 55.6 46.6
StarCoder2-3B Lozhkov et al. (2024) 31.7 27.4 60.2 49.1
CodeGemma-7B-Base Team et al. (2024) 44.5 41.5 65.1 52.4
StarCoder2-7B Lozhkov et al. (2024) 35.4 29.9 54.4 45.6
@@
DeepSeek-Coder-1.3B-Instruct Guo et al. (2024) 65.9 60.4 64.3 54.8
Qwen2.5-1.5B-Instruct Team (2024) 61.6 49.4 63.2 55.6
OpenCoder-1.5B-Instruct Huang et al. (2024) 72.5 67.7 72.7 61.9
Yi-Coder-1.5B-Chat AI et al. (2025) 67.7 63.4 68.0 59.0
Phi-3-Mini-4K-3.8B-Instruct Abdin et al. (2024) 64.6 59.1 65.9 54.2
CodeGemma-7B-Instruct Team et al. (2024) 60.4 51.8 70.4 56.9
@@ 3.1 Experiment I: Performance on Code Generation
We use HumanEval Chen et al. (2021) and MBPP Austin et al. (2021), the most widely used datasets for code generation tasks, as evaluation benchmarks. HumanEval contains 164 tasks, while MBPP includes 500 testing tasks, both equipped with built-in test cases for evaluation. EvalPlus Liu et al. (2023) extends these datasets by introducing stricter test cases to improve assessment robustness. We conduct evaluations using both the original benchmark test cases and their EvalPlus-enhanced versions, denoted by a "+" suffix.
@@
To evaluate the effectiveness of grammar-based code representation, we select DeepSeek-Coder-1.3B Guo et al. (2024) and Qwen2.5-1.5B Team (2024) as baseline models and perform continued pre-training. DeepSeek-Coder-1.3B, trained on high-quality large-scale code datasets, serves as a strong representative of the code model. Meanwhile, Qwen2.5-1.5B-Base, despite being a general-purpose model, demonstrates competitive performance on code-related tasks, making it a valuable point of comparison.
@@
[…]
@@
GrammarCoder models are trained with 8 NVIDIA H800 GPUs. During the base model training phase, we adopt a two-stage learning rate strategy, following approaches from OpenCoder Huang et al. (2024) and MiniCPM Hu et al. (2024). Initially, we use a higher learning rate of 3e-4 to accelerate convergence to a reasonable parameter range. The learning rate is reduced to 5e-5 in the annealing stage for further performance optimization. During the instruct model training phase, we set the learning rate to 5e-5 and trained on an instruction dataset to improve generalization in instruction understanding and code tasks. Throughout the training process, we apply 100 warm-up steps and use a cosine learning rate scheduler to ensure smooth learning rate adjustments, maintaining training stability and efficiency. Additionally, during both token-based and grammar-based continued pre-training, we utilize the same settings to ensure a fair comparison. 2.4 and A.3 provide the detailed information of training dataset.
@@ Results.
Table 1 presents our main experimental results, showing that the GrammarCoder-base model significantly outperforms both the original model and the token-based model trained on the same datasets. For example, on the HumanEval dataset, GrammarCoder-1.3B-Base achieves 82% and 44% improvements over DeepSeek-Coder-1.3B-Base and DeepSeek-Coder-1.3B-Base (CPT), respectively. Notably, after performing continued pre-training on the training dataset, both the token-based and grammar-based models exhibit performance gains. Moreover, even without grammar-based representation, neither the original token-based model nor the continued pre-trained model produces syntax errors, with syntax correctness nearly reaching 100%. Occasional syntax errors (fewer than three) only occur due to random variations on the HumanEval and MBPP datasets. Despite this near-perfect syntax correctness, the grammar-based model still demonstrates superior performance, indicating that incorporating grammar rules provides additional benefits beyond merely preventing syntax errors.
Building on the base model, we further conduct supervised instruction tuning to enhance the model’s adaptability to instruction-based tasks. Table 2 compares the performance of GrammarCoder with current state-of-the-art code models of similar or larger scales. Experimental results show that grammar-based code representation achieves performance comparable to the best token-based models. For example, on the HumanEval (+) dataset, both the base and instruct versions of GrammarCoder outperform other models (e.g., CodeGemma-7B and Yi-Coder-1.5B), while the instruct version achieves performance on par with OpenCoder. However, on the MBPP+ dataset, GrammarCoder-Base does not surpass OpenCoder-Base, which may be attributed to differences in training data volume and quality during the base model pre-training stage. OpenCoder benefits from training on over 900B tokens of high-quality data, whereas GrammarCoder is pre-trained on only around 10B tokens in grammar-based representation. This suggests that while grammar-based representation proves to be effective, the scale and quality of training data also play a crucial role in achieving state-of-the-art performance. Future work can explore expanding the amount of high-quality code data processed into grammar-based representations to further enhance model performance.
@@ 3.2 Experiment II: Understanding the Performance Difference
To validate this hypothesis, we design a new set of experiments focusing on subtle semantic changes that are likely to be overlooked by both humans and token-based models. Specifically, we investigate (1) whether grammar-based representation amplifies these differences, and (2) whether grammar-based models can better capture these changes. These experiments aim to provide deeper insight into how grammar-based representation improves the model’s ability to distinguish between correct and error code, making it more effective in avoiding semantic errors and improving performance.
@@
Qwen2.5-1.5B-Base Team (2024) 72.16 64.97 68.38
@@
We analyze whether grammar-based representation amplifies subtle differences by comparing the edit distances between semantically different code snippets under grammar-based and token-based representations. CodeNanoFix Liang et al. (2024) dataset is used to measure the edit distance, providing a quantitative assessment of how grammar-based and token-based approaches represent code. This dataset consists of 1,000+ programming problems and nearly 100,000 code sample pairs with minimal token differences but significant semantic variations. A subset of 120 programming problems and 3,583 sample pairs serve as the test set. Each sample in the dataset consists of error code submitted by human programmers while solving a problem, along with its corrected version modified by the programmer, both exhibiting minimal token-level differences. Since the differences between error and corrected code typically involve subtle yet crucial semantic changes, such as control flow modifications, variable scope adjustments, and operator usage corrections, this dataset is well-suited for analyzing the differences between token-based and grammar-based representations. To mitigate the impact of outliers, we focus on code pairs with minimal token-level differences (edit distance less than 50, covering 91.18% of the CodeNanoFix’s test set). Additionally, we use GrammarCoder-1.3B’s vocabulary to produce grammar-based representations and DeepSeek-Coder-1.3B’s vocabulary to produce token-based representations, ensuring a fair comparison is made with the maximum overlap of shared tokens.
@@
We evaluate whether grammar-based models more effectively capture these changes by training classifiers using different code representation approaches. Specifically, we use CodeNanoFix as a dataset for a semantic classification task, evaluating the model’s ability to distinguish between semantically correct and incorrect code. By training classifiers to identify code correctness, we examine whether grammar-based representations improve the model’s understanding of code semantics. To ensure a fair comparison, we select baselines that align with GrammarCoder’s architecture. Specifically, we use its corresponding base models, DeepSeek-Coder-1.3B and Qwen2.5-1.5B, along with Condor-1.3B Liang et al. (2024), a model specifically designed for the CodeNanoFix classification task. Precision, Recall, and F1 score are utilized as key metrics for classification performance. Precision evaluates the accuracy of correct code predictions, Recall measures the model’s ability to identify the actual correct code, and F1 score provides a balanced assessment of overall classification performance. Similar to Condor, GrammarCoder is fine-tuned on the CodeNanoFix dataset to enhance its understanding of code semantics and alignment with problem descriptions. In the implementation, a classification layer is added to the original model to output probability scores, with 0.5 sets as the classification threshold. Code snippets with scores above 0.5 are considered correct, while those below are classified as errors. During fine-tuning, the learning rate is set to 5e-5 to ensure stable optimization for the code classification task.

### v2

@@ Abstract
Grammar serves as a cornerstone in programming languages and software engineering, providing frameworks to define the syntactic space and program structure. Existing research demonstrates the effectiveness of grammar-based code representations in small-scale models, showing their ability to reduce syntax errors and enhance performance. However, as language models scale to the billion level or beyond, syntax-level errors become rare, making it unclear whether grammar information still provides performance benefits. To explore this, we develop a series of billion-scale GrammarCoder models, incorporating grammar rules in the code generation process. Experiments on HumanEval (+) and MBPP (+) demonstrate a notable improvement in code generation accuracy. Further analysis shows that grammar-based representations enhance LLMs’ ability to discern subtle code differences, reducing semantic errors caused by minor variations. These findings suggest that grammar-based code representations remain valuable even in billion-scale models, not only by maintaining syntax correctness but also by improving semantic differentiation.
@@ 1 Introduction
Context-free grammars are the fundamental way to specify the syntactic space of a programming language, and with the grammar specified, a program can be parsed into a syntax tree, revealing its structure compiler. Building on this foundation, leveraging grammatical knowledge (e.g., grammar rules) to pre-train large language models (LLMs) has emerged as a promising strategy for code-related tasks, such as code generation grammart5; treegen; graphcodebert2020.
@@
Existing research has explored grammar-based code representation treebert; unixcoder; syncobert; grammart5; treegen; l2s; abstract, where each grammar rule serves as an identity token, and a sequence of grammar rules and terminal tokens represents the program. Figure 1 illustrates a program that determines whether the sum of two integers is odd (top left), along with its corresponding abstract syntax tree (AST) representation (right) and grammar-based representation (bottom left). The grammar-based representation is derived by performing a preorder traversal on the AST. Each grammar rule is extracted independently (e.g., ‘module $\rightarrow$ function_definition’), while terminals are tokenized using a standard tokenizer (e.g., ‘get’). Grammar-based representation has been shown to be effective in preventing syntax errors in encoder-decoder architecture grammart5. Moreover, it facilitates program analysis and enables the pruning of incorrect branches (e.g., filtering out type-error programs l2s; tare) during code generation, thereby enhancing accuracy. Due to these benefits, many code generation models adopt grammar-based representation sun2019grammar; grammart5.
@@
However, as language models scale to the billion-parameter level and beyond, extensive pre-training on large code datasets enables them to implicitly learn syntax rules, making syntax errors increasingly rare chatgpt2024; qwen2; qwen2.5; deepseekcoderv2. For example, even 1B scale models, such as DeepSeek-Coder guo2024deepseekcoder and Qwen2.5 qwen2.5, achieve high accuracy in code generation, consistently producing syntactically valid code. This phenomenon suggests that large models are able to understand the structure of the program and raises a critical question: Is grammar-based code representation still beneficial for billion-scale LLMs?
@@
To answer this question, we conduct an experiment comparing grammar-based representation and token-based representation approaches on 1.3B, 1.5B, and 7B parameter models, respectively. The results demonstrate that grammar-based models (i.e., GrammarCoder-1.3B-Base, GrammarCoder-1.5B-Base, GrammarCoder-7B-Base) significantly outperform token-based models, even though token-based models rarely make syntax errors. For example, on the MBPP dataset mbpp, GrammarCoder-1.3B-Base achieves an almost seven percentage point improvement in Pass@1 compared to DeepSeek-Coder-1.3B-Base trained on the same data. This suggests that grammar rules enhance code generation beyond syntax correction, even in billion-scale models.
@@
  • We release a series of code LLMs trained with grammar rules, providing a valuable resource for further research grammarcoder.
@@ 2.1 Model Overview
We propose GrammarCoder, a grammar-based model built on a decoder-only architecture transformer2017; gpt2018, which excels in auto-regressive tasks like code generation, completion, and translation chatgpt2024; guo2024deepseekcoder; qwen2.5; qwen2.5coder; deepseekcoderv2. To enhance its ability to code generation, we apply continued pre-training and instruction tuning on existing code model weights (i.e., DeepSeek-Coder-1.3B-Base, Qwen2.5-1.5B-Base, and Qwen2.5-7B-Base), expanding its knowledge base. A.1 provides the configuration of the base model we used. In this section, we first introduce our grammar-based code representation. Then, we describe our training strategy and corpus.
@@ 2.4 Training Corpus
We organize our training corpus in two stages: continued pre-training and instruction tuning. Python is selected as the primary programming language for data collection due to its rich syntax and widespread use in diverse programming paradigms. This makes it an ideal candidate for evaluating the effectiveness of grammar-based representations.
For continued pre-training, we sample 10B tokens of Python code from TheStackV2 starcoder2 dataset as the primary training data. Additionally, inspired by previous studies opencoder, we sample 0.5B tokens of self-contained code textbooks from open-source datasets opencoder; nakamura2025aurora to enhance the model’s adaptability to real-world interactive scenarios, bridging the gap between standard pre-training and practical applications.
For instruction tuning, we leverage publicly available instruction datasets opencoder; nakamura2025aurora and employ the data synthesis wei2024selfcodealign; wei2024magicoder approach to collect a total of 6B tokens of instruction data. This ensures the model is better aligned with instruction-following tasks, improving its ability to handle real-world programming scenarios. A.3 provides detailed information about the training datasets.
@@
GrammarCoder-1.5B-Base 65.9 (57.3) 64.8 (55.3)
GrammarCoder-7B-Base 76.8 (71.3) 85.2 (71.7)
Table 1: Comparison of code generation performance between token-based and grammar-based models. The CPT refers to continued pre-training.
@@ 3 Experiments
To evaluate the performance of grammar-based code representations, we develop two sets of models, one with grammar-based code representation and one with token-based code representation. These models are built through continued pre-training from open-source code models, DeepSeek-Coder-1.3B-Base, Qwen2.5-1.5B, and Qwen2.5-7B, on high-quality code data. We begin by evaluating these models on code generation tasks, which are among the most widely recognized and commonly used benchmarks for assessing code-related capabilities (i.e., Experiment I in 3.1).
To further explore the differences between grammar-based and token-based representations, we analyze the reasons contributing to the performance gains of grammar-based representation (i.e., Experiment II in 3.2).
@@
DeepSeek-Coder-1.3B-Base guo2024deepseekcoder 34.8 28.7 56.7 47.9
Qwen2.5-1.5B-Base qwen2.5 37.2 32.9 60.2 49.6
OpenCoder-1.5B-Base opencoder 54.3 49.4 70.6 58.7
Yi-Coder-1.5B yicoder 41.5 32.9 27.0 22.2
CodeGemma-2B-Base codegemmateam2024 26.8 20.7 55.6 46.6
StarCoder2-3B starcoder2 31.7 27.4 60.2 49.1
CodeGemma-7B-Base codegemmateam2024 44.5 41.5 65.1 52.4
StarCoder2-7B starcoder2 35.4 29.9 54.4 45.6
Qwen2.5-7B-Base qwen2.5 57.9 50.6 74.9 62.9
OpenCoder-8B opencoder 66.5 63.4 79.9 70.4
@@
DeepSeek-Coder-1.3B-Instruct guo2024deepseekcoder 65.9 60.4 64.3 54.8
Qwen2.5-1.5B-Instruct qwen2.5 61.6 49.4 63.2 55.6
OpenCoder-1.5B-Instruct opencoder 72.5 67.7 72.7 61.9
Yi-Coder-1.5B-Chat yicoder 67.7 63.4 68.0 59.0
Phi-3-Mini-4K-3.8B-Instruct phi3technicalreporthighly 64.6 59.1 65.9 54.2
CodeGemma-7B-Instruct codegemmateam2024 60.4 51.8 70.4 56.9
@@ 3.1 Experiment I: Performance on Code Generation
We use HumanEval humaneval and MBPP mbpp, the most widely used datasets for code generation tasks, as evaluation benchmarks. HumanEval contains 164 tasks, while MBPP includes 378 testing tasks, both equipped with built-in test cases for evaluation. EvalPlus evalplus extends these datasets by introducing stricter test cases to improve assessment robustness. We conduct evaluations using both the original benchmark test cases and their EvalPlus-enhanced versions, denoted by a “+” suffix.
@@
To evaluate the effectiveness of grammar-based code representation,we select representative and competitive baseline models for comparison, including DeepSeek-Coder-1.3B guo2024deepseekcoder, Qwen2.5-1.5B, and Qwen2.5-7B qwen2.5. DeepSeek-Coder-1.3B, trained on high-quality large-scale code datasets, serves as a strong representative of the code model. Meanwhile, Qwen2.5-1.5B-Base and Qwen2.5-7B-Base, despite being a general-purpose model, demonstrate competitive performance on code-related tasks, making them a valuable point of comparison.
@@
$\text{Pass@1}=\frac{\text{Number of problems solved correctly}}{\text{Total number of problems}}$
@@
GrammarCoder models are trained with 8 NVIDIA H800 GPUs. During the continued pre-training phase, we adopt a two-stage learning rate strategy, following approaches from OpenCoder opencoder and MiniCPM hu2024minicpm. Initially, we use a higher learning rate of 3e-4 to accelerate convergence to a reasonable parameter range. The learning rate is reduced to 5e-5 in the annealing stage for further performance optimization. During the instruction tuning phase, we set the learning rate to 5e-5 and trained on an instruction dataset to improve generalization in instruction understanding and code tasks. Throughout the training process, we apply 100 warm-up steps and use a cosine learning rate scheduler to ensure smooth learning rate adjustments, maintaining training stability and efficiency. Additionally, during both token-based and grammar-based continued pre-training, we utilize the same settings to ensure a fair comparison. 2.4 and A.3 provide the detailed information of training dataset.
@@ Results.
Table 1 presents our main experimental results, showing that the GrammarCoder-Base model significantly outperforms both the original model and the token-based model trained on the same datasets. For example, on the HumanEval dataset, GrammarCoder-1.3B-Base achieves 82% and 44% improvements over DeepSeek-Coder-1.3B-Base and DeepSeek-Coder-1.3B-Base (CPT), respectively. Similarly, on the MBPP dataset, GrammarCoder-7B-Base outperforms DeepSeek-Coder-7B-Base and DeepSeek-Coder-7B-Base (CPT) by 10.3 and 3.7 percentage points, respectively. Notably, after performing continued pre-training on the training dataset, both the token-based and grammar-based models exhibit performance gains. Moreover, even without grammar-based representation, neither the original token-based model nor the continued pre-trained model produces syntax errors, with syntax correctness nearly reaching 100%. Occasional syntax errors (fewer than three) only occur due to random variations on the HumanEval and MBPP datasets. Despite this near-perfect syntax correctness, the grammar-based model still demonstrates superior performance, indicating that incorporating grammar rules provides additional benefits beyond merely preventing syntax errors.
Building on the base model, we further compare our approach with the best-performing models to date and perform instruction tuning to enhance its ability to follow instructions. Table 2 compares the performance of GrammarCoder with current state-of-the-art code models of similar or larger scales. Due to limited computational resources, we perform instruction tuning only on the smaller variants (i.e., GrammarCoder-1.3B-Base and GrammarCoder-1.5B-Base). Experimental results show that grammar-based code representation achieves performance comparable to the best token-based models. For example, on the HumanEval (+) dataset, both the base and instruct versions of GrammarCoder-1.5B outperform other models (e.g., CodeGemma-7B and Yi-Coder-1.5B), while the instruct version achieves performance on par with OpenCoder-1.5B-Instruct. Similarly, on the MBPP (+) dataset, GrammarCoder-7B-Base outperforms the OpenCoder-8B-Base model, further demonstrating the effectiveness of grammar-based representations at larger model scales. However, on the MBPP+ dataset, GrammarCoder-Base-1.5B does not surpass OpenCoder-Base-1.5B, which may be attributed to differences in training data volume and quality during the base model pre-training stage. OpenCoder benefits from training on over 900B tokens of high-quality data, whereas GrammarCoder is pre-trained on only around 10B tokens in grammar-based representation. This suggests that while grammar-based representation proves to be effective, the scale and quality of training data also play a crucial role in achieving state-of-the-art performance. Future work can explore expanding the amount of high-quality code data processed into grammar-based representations to further enhance model performance.
@@ 3.2 Experiment II: Understanding the Performance Difference
To validate this hypothesis, we design a new set of experiments focusing on subtle semantic changes that are likely to be overlooked by both humans and token-based models. Specifically, we investigate (1) whether grammar-based representation amplifies these differences, and (2) whether grammar-based models can better capture these changes. These experiments aim to provide deeper insight into how grammar-based representation improves the model’s ability to distinguish between correct and error code, making it more effective in avoiding semantic errors and improving performance. As we previously observed similar performance trends across different scales of the base models (i.e., results in Table 1), we conduct experiments only at the 1.3B and 1.5B scales in this study to conserve computational resources.
@@
Qwen2.5-1.5B-Base qwen2.5 72.16 64.97 68.38
@@
We analyze whether grammar-based representation amplifies subtle differences by comparing the edit distances between semantically different code snippets under grammar-based and token-based representations. CodeNanoFix liang2024condor dataset is used to measure the edit distance, providing a quantitative assessment of how grammar-based and token-based approaches represent code. This dataset consists of 1,000+ programming problems and nearly 100,000 code sample pairs with minimal token differences but significant semantic variations. A subset of 120 programming problems and 3,583 sample pairs serve as the test set. Each sample in the dataset consists of error code submitted by human programmers while solving a problem, along with its corrected version modified by the programmer, both exhibiting minimal token-level differences. Since the differences between error and corrected code typically involve subtle yet crucial semantic changes, such as control flow modifications, variable scope adjustments, and operator usage corrections, this dataset is well-suited for analyzing the differences between token-based and grammar-based representations. To mitigate the impact of outliers, we focus on code pairs with minimal token-level differences (edit distance less than 50, covering 91.18% of the CodeNanoFix’s test set). Additionally, we use GrammarCoder-1.3B’s vocabulary to produce grammar-based representations and DeepSeek-Coder-1.3B’s vocabulary to produce token-based representations, ensuring a fair comparison is made with the maximum overlap of shared tokens.
@@
We evaluate whether grammar-based models more effectively capture these changes by training classifiers using different code representation approaches. Specifically, we use CodeNanoFix as a dataset for a semantic classification task, evaluating the model’s ability to distinguish between semantically correct and incorrect code. By training classifiers to identify code correctness, we examine whether grammar-based representations improve the model’s understanding of code semantics. To ensure a fair comparison, we select baselines that align with GrammarCoder’s architecture. Specifically, we use its corresponding base models, DeepSeek-Coder-1.3B and Qwen2.5-1.5B, along with Condor-1.3B liang2024condor, a model specifically designed for the CodeNanoFix classification task. Precision, Recall, and F1 score are utilized as key metrics for classification performance. Precision evaluates the accuracy of correct code predictions, Recall measures the model’s ability to identify the actual correct code, and F1 score provides a balanced assessment of overall classification performance. Similar to Condor, GrammarCoder is fine-tuned on the CodeNanoFix dataset to enhance its understanding of code semantics and alignment with problem descriptions. In the implementation, a classification layer is added to the original model to output probability scores, with 0.5 sets as the classification threshold. Code snippets with scores above 0.5 are considered correct, while those below are classified as errors. During fine-tuning, the learning rate is set to 5e-5 to ensure stable optimization for the code classification task.